OpenStudy (anonymous):

Determine whether y varies directly with x. If so, find the constant of variation k and write the equation.

OpenStudy (anonymous):

so whats your variables?

OpenStudy (anonymous):

I can give you some for an example and show you?

OpenStudy (unklerhaukus):

\[y\propto x\]\[y=kx\]

OpenStudy (anonymous):

like -6y = -5x If y varies directly as x, then y = kx. -6y = -5x y = -5x/-6 y = 5x/6 y = (5/6) * x

OpenStudy (anonymous):

So, yes, y does vary directly with x, and k = 5/6.

OpenStudy (anonymous):

The graph of this equation is a straight line passing through the origin with gradient = k = 5/6. This is if you were looking for a graph